calculate the number of moles of hydrogen gas present in 0.5 cm cube sample of hydrogen gas at pressure of 101.325 KP aaa and aa
temperature 27 degree Celsius

Grade:11

1 Answers

Arun
4 years ago
The data given in the problem isP = 760mm of Hg = 1 atm V = 500cm3 T =  270C T = 27+ 273 = 300 K R = 82.1 cm3 K-1mol-1 number of moles of hydrogen (n) = ? According to ideal gas equation PV= nRT n = PV/RT = 1 atm X 500 cm3 /82.1 cm3 K-1mol-1 X 300 K n = 0.023mol